Specifically the price of tickets between Lusaka and Livingston’s increased from 200 kwacha to 250 kwacha . The demand equation is Qd =200-5p. Calculate the old price quantity demanded and the new price quantity demanded level


1
Expert's answer
2021-09-07T08:04:01-0400

"Qd = 200 - 5( 200)"


"= 200-1000 =- 800"

"New Qd = 200-250(5)"

"= 200-1250 = 1050"